OpenStudy (anonymous):

Helpp. Solve the equation –5 × 1 × 11 × 4 step by step. List each step, along with the property you used during that step.

OpenStudy (tkhunny):

It cannot be done. It's not an equation. The expression can be evaluated, however. Use your Order of Operations and start on the left.

OpenStudy (anonymous):

-5*11*4=-55*4=-220

OpenStudy (anonymous):

is -220